Solve for X to the nearest thousandth:
100 ln(2X-3)- 10 = 102
I think I should add 10 to both sides of the equation so it looks like:
100 ln(2x-3) = 112.
Am I right so far? Now I'm stuck :(

Answer by lwsshak3(11628)

100 ln(2X-3)- 10 = 102
..
100 ln(2X-3) = 102+ 10
ln(2X-3) = 112/100=1.12
convert to exponential form
e^1.12=2x-3
using calculator
3.06485..=2x-3
2x=6.065
x=3.033 (ans)
You were on the right track. You just needed a little more work to finish.